OpenStudy (anonymous):

which of the following is equal is the rational expression when x= 5 or -1 (x-7)(x+1)/(x+1)(x-5)

OpenStudy (anonymous):

\[x \neq5\]

OpenStudy (noelgreco):

The question is worded oddly.

OpenStudy (anonymous):

yeah i know its an multiple choice question

OpenStudy (anonymous):

@NoelGreco

OpenStudy (noelgreco):

That's not what I mean. The words before the math expression don't make sense.

OpenStudy (xapproachesinfinity):

rewrite your problem! it is not written neatly or just take a snapshot and attach it

OpenStudy (anonymous):

which of the following is equal to the rational expression when \[x \neq5 or -1\]

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (anonymous):



OpenStudy (xapproachesinfinity):

the question is inadequate! you are missing info

OpenStudy (xapproachesinfinity):

you better take a picture of the question and post it and we will see

OpenStudy (xapproachesinfinity):

all i can tell you now is that if \(\huge \rm x\ne-1,5\) then \(\huge \rm \frac{(x-7)(x+1)}{(x+1)(x-5)}=\frac{(x-7)}{(x-5)}\)

OpenStudy (anonymous):

thats what i got to, thats all they wanted to know i thought it was more thats why i asked for the help my bad for the confusing

OpenStudy (xapproachesinfinity):

make sure to port the whole question next time lol